OpenStudy (anonymous):

Two automobiles leave the same location traveling in opposite directions. The faster car travels 55 mph and leaves one hour before the slower one, which travels 47 mph. Use x for the time traveled by the faster car and y for the time traveled by the slower car. Write the equations necessary to find how long before the are 350 miles apart. Do not solve.

OpenStudy (kropot72):

x = y + 1 ................(1) The separation speed when both cars are travelling is 55 + 47 = 102 mph The distance travelled by the fastern the initail one hour = 55 miles, therefore the remaining distance to be travelled at the separation speed of 102 mph is 350 - 55 = 295 miles. So the time taken by the slower car is y = 295/102 hours ..........(2) Substituting the value for y obtained in (2) into equation (1) the time taken for the cars to be 350 miles apart is the time x which is given by x = (295/102) + 1

OpenStudy (kropot72):

the faster car in the initial*

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

You're welcome :)